how do you slove in combination/elimination method :
-4x - y = -6
4x + 3y = 18

Answer by jim_thompson5910(35256)

Add the equations to get


-4x - y = -6
4x + 3y = 18
-------------
0x + 2y = 12


So, 2y = 12 ----> y = 6 (divide both sides by 2)


Now that we know that y = 6, we use this to find x.


-4x - y = -6

-4x - 6 = -6

-4x = -6+6

-4x = 0

x = 0/(-4)

x = 0


So the solutions are: x = 0, y = 6